OpenStudy (shamil98):

f(x) = x^3 × arccot(4x) Just need to check my work..

OpenStudy (shamil98):

Using the product and chain rules.. f'(x) = x^3 × ( arccot(4x))' × (x^4) × (arccot(4x))

OpenStudy (shamil98):

times (x^3) my bad.. f'(x) = x^3 × - 1/(1+x^2) × (4x)' × 3x^2 × arccot(4x)

OpenStudy (shamil98):

f'(x) = x^3 × -1/(1+x^2) x 4 x 3x^2 x arccot(4x)

OpenStudy (shamil98):

\[f'(x) =x^3 × \frac{ -1 }{ 1 + x^2 } \times 4 \times 3x^2 \times arccot(4x)\]

OpenStudy (anonymous):

f(x) = x^3 × arccot(4x) f '(x) = x^3 X -1/(1+(4x)^2) + 3x^2 X arccot(4x) simplify and answer :)

OpenStudy (shamil98):

oh oops forgot about the 4x in the denominator.

OpenStudy (anonymous):

the only thing you are doing wrong is you are not replacing the x in -1/(1+x^2) by 4x which is in arccot(4x)

OpenStudy (shamil98):

yeah, thanks .

OpenStudy (anonymous):

you are welcome, if you can help me on my question lol